Quiz: Quiz Chapter 2 Time Limit: 02 00.00 Submit Quz This Question: 1 pt 7 of 15 (7 complete) This Quiz: 15 pts possible The supply of coffe is Qs-2.0+0.2P where quantity is in billions of pounds and price is dollars per pound. The price elasšicly of demand for coffe is about -0 3, the current price is $3.00 per pound and the quantity bought and sold is 2 6 bilion pounds Based on this information, the linear demand curve for coffeeis Q term, so enter your slope coefficient as a positive number) A severe drought reduces the supply of coffee by 10 percent at every price. The new equililbriurm price of coffee will be and the new equilbrium quantity will be (Enter your responses-the intercept and slope - as real numbers rounded to 2 decimal places Note that a negative sign already appears in front of the slope billion pounds. (Enter your responses as real numbers rounded to 2 decimal places)
Answer #1

The linear demand curve is of the form, Q=a-bP, where b represents the elasticity value

2.6=a-0.3*3

a= 3.5

So, the linear demand curve is Qd = 3.5-0.3P

The new equilibrium price of coffee will be $3.54 and the new equilibrium quantity will be 2.44 billion pounds

P

Qd

Qs

New Qs

0

3.50

2

1.8

0.5

3.35

2.1

1.89

1

3.20

2.2

1.98

1.5

3.05

2.3

2.07

2

2.90

2.4

2.16

2.5

2.75

2.5

2.25

3

2.60

2.6

2.34

3.54

2.44

2.708

2.44

4

2.30

2.8

2.52

4.5

2.15

2.9

2.61